# unPIC: A Geometric Multiview Prior for Image to 3D Synthesis

### v2

@@ 4.1 What features are most useful to decode novel views?
To motivate geometric representations and the use of CROCS specifically, we run an experiment to decode a single novel view of an object at a fixed rotation ($\theta=90$ degrees). We train diffusion models that take a source image and a ground-truth annotation of the target image to predict the target image. The annotations range from geometric (like depth maps and CROCS), image-space (alpha masks), through feature maps (from DINOv2 [35] and CLIP [38]). All annotations are treated as images—they are upscaled (if required) and concatenated to the input latent before it is fed to the UNet for denoising. We run these experiments at 64x64 image size to allow multiple seeds.
We find in Table 1 that geometric annotations consistently outperform other choices. Given a source image, there is more ambiguity about the target shape than its texture or appearance, explaining the empirical results. We also see a benefit of CROCS over NOCS: coloring the annotation consistently makes it more useful than feeding in arbitrary colors. The full benefit of CROCS will be revealed when we try to predict CROCS and NOCS each in Sec 4.2.
Multiple targets. To examine the benefit of geometric features for multiview prediction, we run experiments with and without ground-truth annotations when predicting 4 or 8 target views. The target superimage dimensions are 128x128 and 256x128, respectively. We carefully control the diffusion architecture, applying attention at two downsampled resolutions in each case (when the feature height dimension is 16 or 8).
In Table 2, we see a large improvement in the MSE by conditioning on multiview CROCS pointmaps. In fact, we see increasing benefits from $K=1$ (using CROCS is $3.9$x better) to $K=4$ ($4.5$x better) and $K=8$ ($5.6$x better). The CROCS-conditional error is not only smaller relative to the unconditional case, it actually decreases in absolute terms when predicting more views. This is remarkable because the number of tokens doubles with the number of views, making it harder to model spatial relationships across the whole superimage (and perhaps leading to the drop in the unconditional case). These results help establish the advantage of (ground-truth) CROCS annotations in coordinating multiview synthesis.
Table 1: Single-target diffusion decoder experiments. We predict a fixed 90-degree object rotation from a given source image, with the help of a ground-truth annotation of the target image. We report the pixel-space Mean Square Error on random source views across 10k heldout objects. Stds. are across 3 training runs.
@@
Table 2: Multiple-target diffusion decoder experiments using ground-truth annotations. We report the pixel MSE (1e-3) on $K-1$ novel views as in Table 1. Stds. are across 5 training runs.
Annotation Type K=4 views K=8 views
Source image only 4.739 $\pm$ 0.215 5.817 $\pm$ 0.230
+ CROCS (K views) 1.065 $\pm$ 0.064 1.039 $\pm$ 1.331
4.2 What multiview features are predictable from a single image?
Having used ground-truth representations in Section 4.1, we now move to estimating multiview geometric representations from a single source image. We train a diffusion prior to predict pointmaps at 8 target views. We compare two choices of representations for how predictable they are from arbitrary source views: (i) NOCS images in a static reference frame, and (ii) CROCS, where we change the reference frame based on the source camera.
We find a 4x increase in predictability from using CROCS (see Table 3). Without CROCS, choosing a random source camera leaves the colors of the target images unpredictable because we do not condition the model on any camera extrinsics. With CROCS, each target pose has a biased distribution of colors, consistent across objects. This makes the geometric feature space unambiguous, and each representation loop more predictable, in turn freeing up the model’s probabilistic capacity for the ambiguous task of predicting unseen object/scene geometry.
Table 3: Diffusion prior experiments. We predict pointmaps at 4 and 8 target views, and report the pointmap MSE (1e-3).
Annotation Type K=4 views K=8 views
NOCS 11.94 15.82
CROCS 1.21 3.92
Table 4: Comparison with baselines on 4 datasets.
Dataset Method PSNR $\uparrow$ IoU $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ CLIP MV $\downarrow$
Objaverse XL CAT3D 22 0.63 84 0.23 0.75 2.44
EscherNet 12 0.62 79 0.26 0.65 2.36
@@ 4.2 What multiview features are predictable from a single image?
One-2-3-45 12 0.58 68 0.25 0.66 2.55
OpenLRM 14 0.62 103 0.23 0.71 2.38
unPIC-latents (ablation) 23 0.71 78 0.18 0.79 2.23
unPIC-pixels (ours) 25 0.79 83 0.19 0.78 2.16
Google Scanned Objects CAT3D 24 0.68 78 0.23 0.75 3.08
EscherNet 14 0.73 72 0.25 0.67 2.79
Free3D 16 0.73 83 0.23 0.71 3.60
One-2-3-45 13 0.65 57 0.25 0.66 2.74
OpenLRM 14 0.67 101 0.25 0.69 2.93
unPIC-latents (ablation) 23 0.80 69 0.18 0.78 2.42
unPIC-pixels (ours) 26 0.89 87 0.18 0.79 2.63
Amazon Berkeley Objects CAT3D 22 0.64 70 0.26 0.71 2.37
EscherNet 11 0.65 70 0.30 0.60 2.24
Free3D 14 0.73 64 0.24 0.66 2.87
One-2-3-45 11 0.61 48 0.27 0.62 2.22
OpenLRM 12 0.59 115 0.29 0.65 2.55
unPIC-latents (ablation) 25 0.83 50 0.16 0.78 1.88
unPIC-pixels (ours) 27 0.87 58 0.17 0.78 1.74
Digital Twin Catalog CAT3D 23 0.74 68 0.20 0.76 2.70
EscherNet 14 0.75 71 0.22 0.69 2.29
Free3D 17 0.80 75 0.19 0.72 2.86
One-2-3-45 14 0.74 49 0.20 0.70 1.98
OpenLRM 14 0.71 88 0.22 0.72 2.40
unPIC-latents (ablation) 25 0.84 60 0.14 0.82 1.89
unPIC-pixels (ours) 27 0.91 68 0.16 0.81 1.81
5 Main Results
We compare unPIC on single-image-to-multiview synthesis with the following state-of-the-art methods:
  1. 1.
  2. 2.
    EscherNet (256x256): a geometry-free diffusion model equipped with a specialized camera encoding (CaPE), allowing a flexible number of input images and target views.
  3. 3.
  4. 4.
  5. 5.
Figure 4: Qualitative comparison. The left column shows the source image, while the remaining images in each row are predicted novel views. One-2-3-45 produces multiview inconsistencies. CAT3D can squash the shapes in unseen views.
Our model is based on pixel-diffusion and trained from scratch on synthetic data only. In contrast, most competing methods have the advantage of starting from a pretrained model: Free3D, EscherNet, and Zero-1-to-3 start from Stable Diffusion; CAT3D starts from a similar pre-trained latent diffusion model [40]; and OpenLRM bootstraps from DINO features. The large-scale pretraining of these models/features is an obvious advantage.
An ablation. To demonstrate that training from scratch on synthetic data is not necessarily a disadvantage, we run the following baseline: we train a latent diffusion model (denoted unPIC-latents) on the same training data as our main model (denoted unPIC-pixels in Table 4 or simply unPIC elsewhere). We use the VAE from Stable Diffusion 1.4 to encode all CROCS pointmaps, and source and target images. The VAE compresses images from 256x256x3 down to 32x32x4. We then train a diffusion model maintaining a similar architecture to unPIC-pixels. We concatenate latents for $K=8$ views to form a 4x2 grid as we did in the image case. Although the dimensionality of the latent grid is smaller, we keep all architectural hyperparameters the same, allowing unPIC-latents a slight advantage.
Test datasets. In addition to our 50k held-out examples of Objaverse XL, we evaluate on three out-of-distribution datasets: (i) Google Scanned Objects [8], (ii) Amazon Berkeley Objects [4], and (iii) the Digital Twin Catalog [36]. Some of the baselines may have trained on these test datasets—for instance, One-2-3-45 was trained on ObjaverseXL (we use the standard Zero123-XL checkpoint).
  • Geometric accuracy: we compute the IoU (Intersection over Union) of the 2D mask of a predicted novel view w.r.t. the expected mask. This metric is invariant to lighting or texture.
  • Multiview consistency: we compute CLIP embeddings (using the open-source ViT L/14 336px model) for all K generated images in a given example, then compute the mean of the (K x K) pairwise distance matrix of the embeddings. This metric (abbreviated CLIP-MV) avoids any comparison with ground-truth images, and merely probes the internal consistency of a set of images. All reported CLIP-MV numbers are 1e-4.
When computing metrics, we sample a random target from the set of 7 novel views for each example (except for CLIP-MV, where all novel views are used). This ensures the computed statistics are independent (particularly because FID relies on second-order statistics). We do not repeat any object (as in training) with an alternative source view. We run all models, including ours, on A100 GPUs, and downsize all images to 256x256 for comparison.
Results. Table 4 shows that unPIC beats all baselines on PSNR and SSIM. Notably it outperforms them substantially on the shape-focused IoU metric. We examine this advantage qualitatively in Figure 4. Although all methods tend to produce plausible looking images, the generated poses and shapes can be completely arbitrary (Figure 5). Being trained on Objaverse and ObjaverseXL alone, unPIC underperforms on the FID score in comparison to methods which likely benefit from ImageNet pretraining. Nevertheless, its LPIPS scores establish the perceptual quality of its outputs.
Figure 5: While all models produce plausible images, their shapes and poses can be off. We compare masks with the ground-truth.
We find a common failure mode that affects shape and pose prediction for CAT3D and OpenLRM: despite specifying a constant camera height, OpenLRM appears to level and raise the camera (w.r.t. the ground plane) as it moves around the object (see fig. 2 in the LRM paper [18]). The issue appears consistently in OpenLRM outputs, and intermittently in CAT3D’s as well. It likely stems from how the training data was generated for those models: the orientation of each object was likely reset before capturing images. As a result, CAT3D ends up modeling the stochasticity of the object’s initial orientation; OpenLRM, being a deterministic model, overfits to it. At inference time, these methods randomly guess an orientation, so they may assume the object is tilted w.r.t. the world’s up-direction. This leads to an apparent camera-control failure, where the camera appears to raise or lower itself as it moves around the object. It is especially undesirable for assets placed on a planar surface, as the whole surface appears to tilt while the camera orbits the object. In our case, we avoid this issue by ensuring the object’s vertical axis is always aligned with the world’s.
@@ 6 Discussion
Figure 6: Diversity of unPIC outputs from an ambiguous image (top-left). All samples show a 90-degree rotation from the source view, and are shown here without cherry-picking.
Figure 7: In-the-wild inference on natural images from two datasets: GQA [20] and Objectron [1]. As in Figure 1, the first column shows the source image; columns 2-5 show four of eight predicted CROCS pointmaps; and columns 6-9 show corresponding novel-view predictions. unPIC generalizes well beyond the digital assets it was trained on.
Multiview synthesis is central to image-to-3D modeling. A sparse but targeted set of novel views can serve as an ”anchor” for downstream generation of more fine-grained views, an arbitrary target pose, or an explicit 3D reconstruction. This underscores the importance of improving existing multiview priors, which tend to be image-focused and geometry-free.
There are several advantages of our hierarchical approach: first, it allows maximizing the size of the prior and decoder models, as they can be trained separately; second, we can visualize and interpret the intermediate features—for instance, visualizing CROCS could reveal whether 3D reconstruction errors arise due to geometry or texturing; and finally it prevents a collapse of output diversity in the prior—this would be much likelier if the modules were cotrained, as the prior would be spurred to output only the optimal intermediate representation for the final RGB loss. See Figure 6 for an example of diverse 3D shapes predicted from a single image.
We posit that unPIC’s main advantage is being geometry-driven. This was evidenced by the geometric accuracy and multiview consistency of unPIC’s novel-view predictions in Section 5. To provide further evidence that decoupling geometry and appearance supports generalization, we run unPIC on in-the-wild images far from its training distribution (see Figure 7). These results suggest that unPIC could be useful for a variety of downstream 3D applications, and help ensure that downstream tasks are well-posed.
Limitations & Future Work. A salient shortcoming is that we do not model scene backgrounds. This may be feasible using CROCS for room-like scenes where the background geometry is uniform, and can be scaled to a unit cube. We are currently also limited to training on images of 3D assets rendered at controllable poses, rather than in-the-wild images. This could be eased by using pretrained 3D models (e.g., NeRFs) of real scenes to render images and depth maps, which then could be used to estimate CROCS.
A straightforward extension of our study would be to allow multiple random source views rather than one. One of them could be designated the primary view. It would be interesting to compare the model’s performance in two cases: when it infers the correspondence between the source images, or when they are accompanied by camera pose annotations to avoid correspondence ambiguity.
While our ablation comparing the latent-diffusion implementation of unPIC with the pixel-diffusion-based implementation did not suggest any immediate advantage for unPIC-latents, this was likely due to the fact that the pretrained VAE did not offer a suitable latent space to encode CROCS pointmaps. It would be worth fine-tuning a VAE or training a geometry-focussed latent space.
The unPIC framework is not limited to one prior, e.g., it could be fruitful to predict multiview texture features to condition the decoder. It may also be possible to enhance multi-object handling by breaking CROCS into two components: a scene-level pointmap like ours (that colors objects based on their positions in the scene), and a pointmap that colors objects individually, capturing each object’s shape independent of its position in the scene.
Conclusion. We introduced the unPIC framework for multiview synthesis from a single image via multiview features. Our choice of intermediate features, CROCS, encodes scale-free geometry and point-to-point correspondence across views. We showed it is an effective choice over several geometric and non-geometric alternatives. It is predictable irrespective of the camera pose of the source image, and supports learning and generalization across examples.
Our two-tier system helps address the under-specification of image-to-3D tasks by allowing probabilistic outputs at each level. We encourage multiview consistency at each level using explicit geometric supervision and control. Our approach outperforms multiple competing alternatives on a variety of metrics, especially in terms of shape and pose accuracy. unPIC learns shape priors that generalize well beyond the digital assets it was trained on, paving the way for more geometry-driven novel view synthesis.
Appendix A Layout
The appendix is organized as follows: we describe CROCS in detail in Appendix B, discussing in particular the geometrical properties which led to our algorithmic choices. We then present additional qualitative comparisons and quantitative analyses in Appendix C. We break down unPIC’s hierarchical error in Section C.2. Finally, we describe all training and evaluation details for unPIC in Appendix D.
Appendix B Geometric Subtleties
B.1 CROCS Rescaling

### v3

@@ 4.1 Choice of Intermediate Representations
Predictiveness. We train a primitive version of unPIC’s decoder module to predict a single novel view of an object at a fixed target rotation ($\theta=90$ degrees). The model is conditioned on a source image, as well as a ground-truth annotation of the target novel view. The type of annotation is varied across training runs to assess the usefulness of different representations. The annotations range from geometric (like depth maps and CROCS), image-space maps (alpha masks), through feature maps (like DINOv2 [30] and CLIP [33]). All annotations are treated as images—they are upscaled (if required) and concatenated to the input image before it is denoised. We run these experiments at 64x64 image size (to feasibly upscale the high-dimensional CLIP and DINOv2 features), and using pixel-based diffusion to avoid ambiguity about how to encode the annotations. Table 1(a) shows that CROCS is the most helpful for this task.
Table 1: We analyze the effectiveness of CROCS for novel view synthesis (top) and its predictability from a single image (bottom) versus other geometric representations.
(a) Does CROCS help predict novel views? We predict a $90^{\circ}$ object rotation using varied target image annotations. We report pixel MSE and stds. over 3 training runs.
@@
(b) Which multiview features are predictable from a single image? We report the pointmap Mean Square Error (1e-3) at K=4 and K=8 target views.
K Annotation Type MSE (1e-3)
4 NOCS 11.94
CROCS 1.21
8 NOCS 15.82
CROCS 3.92
Predictability. Having shown that CROCS improves the decoder, we now ask: can CROCS itself be predicted reliably? We train primitive versions of the unPIC prior to predict either NOCS or CROCS from a single source image. We experiment with $K=4$ or $K=8$ target views. We compare two choices of representations: (i) NOCS images in a static reference frame, determined by the default creator-intended pose for the given asset; and (ii) CROCS, where we change the reference frame based on the source camera. We find a 4x increase in predictability from using CROCS (see Table 1(b)). With CROCS, each target pose has a biased distribution of colors, consistent across objects. This makes the geometric feature space unambiguous, freeing up the model’s probabilistic capacity for the ambiguous task of predicting the unseen object/scene geometry.
4.2 Task 1: Novel View Synthesis
We compare unPIC against a diverse set of seminal and state-of-the-art baselines: CAT3D, EscherNet, Free3D, OpenLRM, InstantMesh (Large), and One-2-3-45 XL. The key conceptual differences are summarized in Table 2 and detailed further in Appendix B.7.
Test data. We sample 512 test objects from 4 different datasets, including our held-out subset of Objaverse XL, and three out-of-distribution datasets: (i) Google Scanned Objects [6], (ii) Amazon Berkeley Objects [3], and (iii) the Digital Twin Catalog [31]. The baselines may have trained on these test datasets—for instance, One-2-3-45 was trained on ObjaverseXL. For each test object, we sampled a random source and target view, and evaluated all models on the same pair.
  • 2D Geometric Accuracy: We compute the IoU (Intersection over Union) of the binarized mask of the predicted novel view w.r.t. the expected mask. This metric is invariant to lighting or texture.
  • Multiview Consistency: Though we attempted to run MET3R [1] to measure multiview consistency, we found it inadequate for our use case, even for ground-truth images (see Appendix B.6). Instead, we compute CLIP (ViT L/14 336px) embeddings for K=8 generated images per object, including the source image. Then, we compute the mean of the (K x K) pairwise distances between the embeddings. This metric (abbreviated CLIP MV) helps probes the consistency of a set of generated images without comparing to ground-truth images.
Results. unPIC consistently outperforms the baselines on all novel-view metrics (Table 3). The IoU metric provides an early sign of its geometric accuracy. On the CLIP MV metric, it is noteworthy that the only method which outperforms unPIC consistently, InstantMesh, is also geometry-supervised.
Table 2: Summary of different approaches for single-image-to-3D generation. P: Probabilistic/Generative; D: Deterministic; $F$ Recon: Feed-Forward Reconstruction; V-Gen $\rightarrow$ Opt: View Generation followed by per-scene Optimization; SVD: Single-View Diffusion (run repeatedly per target viewpoint); MVD: Multi-View Diffusion (generates views jointly).
Model Output Model Type Generative Core Intermediate 3D Rep. Direct Geometric Supervision
LRM D FF Recon - Triplane $\rightarrow$ NeRF No (Image loss only)
Free3D P V-Gen $\rightarrow$ Opt MVD Implicit (No 3D Rep) No (Image loss only)
CAT3D P V-Gen $\rightarrow$ Opt MVD Implicit (3D Attention) No (Image/Perceptual loss)
EscherNet P V-Gen $\rightarrow$ Opt MVD Implicit (MV Attention) No (Image loss only)
One-2-3-45 P V-Gen $\rightarrow$ FF Recon SVD SDF (from Cost Vol.) Depth
InstantMesh P V-Gen $\rightarrow$ FF Recon MVD Triplane $\rightarrow$ Mesh (FlexiCubes) Depth & Normals
unPIC (ours) P Direct 3D + V-Gen MVD (Hierarchical) Pointmaps (CROCS) Pointmaps (CROCS)
Table 3: Novel-View Synthesis: empirical comparison on 4 datasets.
Dataset Method PSNR $\uparrow$ IoU $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ CLIP MV (1e-4) $\downarrow$
Objaverse XL CAT3D 12 0.61 82 0.27 0.64 2.52
EscherNet 13 0.63 87 0.24 0.68 3.13
InstantMesh 13 0.62 76 0.25 0.67 1.89
@@ 4.2 Task 1: Novel View Synthesis
One-2-3-45 12 0.57 72 0.26 0.65 2.55
OpenLRM 13 0.60 94 0.24 0.71 2.38
unPIC (ours) 24 0.72 63 0.15 0.80 2.01
Google Scanned Objects CAT3D 19 0.64 87 0.26 0.72 2.98
EscherNet 14 0.72 76 0.26 0.66 2.78
InstantMesh 14 0.71 61 0.25 0.66 2.12
Free3D 14 0.70 93 0.25 0.68 3.63
One-2-3-45 13 0.64 60 0.26 0.65 2.74
OpenLRM 14 0.66 109 0.26 0.69 2.93
unPIC (ours) 23 0.77 62 0.17 0.77 2.33
Amazon Berkeley Objects CAT3D 18 0.59 82 0.29 0.67 2.34
EscherNet 11 0.65 73 0.30 0.59 2.35
InstantMesh 12 0.67 60 0.27 0.62 1.54
Free3D 12 0.69 73 0.27 0.63 2.92
One-2-3-45 11 0.61 50 0.28 0.61 2.22
OpenLRM 11 0.57 122 0.30 0.65 2.55
unPIC (ours) 25 0.88 40 0.15 0.78 1.74
Digital Twin Catalog CAT3D 18 0.70 83 0.23 0.71 2.68
EscherNet 14 0.74 75 0.23 0.68 2.37
Free3D 14 0.76 87 0.22 0.70 2.90
InstantMesh 14 0.74 58 0.22 0.68 1.51
One-2-3-45 14 0.74 53 0.21 0.69 1.98
OpenLRM 14 0.70 98 0.22 0.72 2.40
unPIC (ours) 25 0.90 46 0.13 0.82 1.55
Table 4: Ablation: we train a non-hierarchical version of unPIC without any geometry.
Dataset Method PSNR $\uparrow$ IoU $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ CLIP MV (1e-4) $\downarrow$
Objaverse XL unPIC ablation 20 0.63 89 0.23 0.75 2.07
Google Scanned Objects unPIC ablation 21 0.69 97 0.24 0.75 2.43
Amazon Berkeley Objects unPIC ablation 19 0.61 84 0.28 0.70 1.84
Digital Twin Catalog unPIC ablation 22 0.82 69 0.19 0.78 1.67
Table 5: 3D Geometric Accuracy: we compute the bidirectional Chamfer distance (1e-3) $\downarrow$ to compare point cloud completeness and accuracy. All methods take a single image as input. GSO: Google Scanned Objects. DTC: Digital Twin Catalog.
Method GSO DTC
EscherNet [21] 31.4 -
Direct3D [49] 27.1 -
Zero123++ & MeshLRM [47] 15.7 -
Direct3D (re-run) 8.94 5.99
InstantMesh (re-run) 6.83 6.62
unPIC (ours) 4.59 5.47
Figure 3: Novel views, baselines vs unPIC. The left column shows the source view, as output by each model. The models show different failure modes: (i) mutations of the original object (e.g., InstantMesh, EscherNet; annotated in blue); (ii) multi-view inconsistencies (e.g., One-2-3-45, Free3D; annotated in red); (iii) issues inferring the depth of the boat from the source image (e.g., CAT3D, OpenLRM); and (iv) pose inaccuracies in the novel views (e.g., InstantMesh gets the pose of the sail right, but the pose of the boat wrong; annotated in violet).
Figure 4: Points clouds, baselines vs unPIC.
4.3 Task 2: 3D Reconstruction
We derive point clouds from unPIC by simply combining the multiview outputs. The CROCS images provide the vertices, while the RGB images provide the vertex colors; we only need to mask out the background pixels from the output images. We run two competing methods, Direct3D and InstantMesh (Large), on our test images for comparison. To ensure the Chamfer distances are comparable, we subsample 8192 points from every point cloud. We also provide the numbers reported by EscherNet, MeshLRM, and Direct3D; our Direct3D re-run produced a similar score as the original paper. Table 5 shows that unPIC outperforms all methods. See Figure 4 for a qualitative comparison.
4.4 Ablating Hierarchical Generation
We train a version of unPIC without any geometry inputs or supervision, under identical training conditions. It is non-hierarchical as it takes a single RGB image and outputs the target novel views. In Table 4, we find that the ablated model performs worse than the full version of unPIC; nevertheless, it is comparable to CAT3D, which is another geometry-free multiview synthesis method.
@@ 5 Discussion
Figure 5: Diversity of outputs. Top row: various test images from DTC and ObjXL. Remaining rows: unPIC’s output at a fixed $90^{\circ}$ rotation across three RNG seeds, showing the range of outputs for the given conditioning.
Generative diversity. We qualitatively demonstrate the advantage of a hierarchical approach by sampling multiple outputs for a fixed input image. Figure 5 shows that unPIC’s novel views remain plausibly constrained while exhibiting appropriate variety across random seeds. A hierarchical pipeline like unPIC helps capture multimodal distributions in both geometry (varying structural interpretations of ambiguous inputs, e.g., the pomegranate or toy blocks) and appearance (texture variations, e.g., the white device).
Generalization and real-world utility. Having never seen a real-world pixel during its training, unPIC nevertheless generalizes to in-the-wild images (see Figure 23). This is clear evidence that it is possible to learn real-world shape priors from synthetic assets alone. While our focus was on using these priors for image-to-3D generation, unPIC could be repurposed for vision tasks, e.g., monocular or stereo depth estimation.
Limitations and failure cases. (i) A salient shortcoming is that we do not model image backgrounds. This may be feasible while still using CROCS where the background geometry is uniform, and can be scaled to a unit cube (e.g., room-like scenes). (ii) unPIC rotates some objects in the image plane rather than spinning them around the vertical axis, likely because it misinterprets the source image as an overhead view of the object. Since CROCS does not account for the source camera elevation, we could potentially condition unPIC on the elevation to prevent ambiguity. We provide an example of this failure in Figure 24. (iii) Natural images containing faces or multiple people can be out-of-distribution, as unPIC was trained on synthetic assets lacking such cases. We hope that expanding our training data, e.g. by combining multiple assets in the same scene, could offer one avenue to ease this limitation. (iv) At present, unPIC is also limited by the expressiveness of the SD-VAE: for example, it cannot represent any text. A richer representation space could help get over this limitation.
Conclusion. We introduced the unPIC framework for image-to-3D generation. Our choice of intermediate features, CROCS, encodes point-to-point correspondence across views in a canonical reference frame. CROCS allows us to invert the usual “generate then reconstruct” paradigm, by producing 3D directly as part of the view-generation process. With its hierarchical approach and explicit geometric supervision, unPIC outperforms numerous competing methods on novel-view synthesis, multiview-generation consistency, and 3D geometric accuracy. unPIC learns shape priors that generalize well beyond the digital assets it was trained on, paving the way for more geometry-driven novel-view synthesis.
Appendix A Appendix
A.1 CROCS
We first provide a precise specification of CROCS.
Normalization (similar to NOCS)
Let $\Omega\subset\mathbb{R}^{3}$ denote the surface of a 3D object or scene. For any point $\mathbf{p}=(x,y,z)\in\Omega$ in world coordinates, we first normalize its coordinates into NOCS, defined by the unit cube $0,1]^{3}$.
We determine the Axis-Aligned Bounding Box (AABB) of $\Omega$ using the minimum and maximum coordinates, $\mathbf{p}_{\text{min}}$ and $\mathbf{p}_{\text{max}}$. The geometric center of the AABB is $\mathbf{c}_{\Omega}=(\mathbf{p}_{\text{min}}+\mathbf{p}_{\text{max}})/2$. To ensure uniform scaling while preserving the aspect ratio, we calculate the maximum extent $L$ of the bounding box:
$L$ $=\|\mathbf{p}_{\text{max}}-\mathbf{p}_{\text{min}}\|_{\infty}$ (1)
$=\max(x_{max}-x_{min},y_{max}-y_{min},z_{max}-z_{min}).$ (2)
The NOCS coordinates $\mathbf{p}^{\prime}\in[0,1]^{3}$ are then obtained by scaling the object and centering it at $\mathbf{h}=(0.5,0.5,0.5)$:
$\mathbf{p}^{\prime}=\frac{1}{L}(\mathbf{p}-\mathbf{c}_{\Omega})+\mathbf{h}.$ (3)
Camera-Relative Transformation
Let the position of a source camera be defined in spherical coordinates by $(\theta,\phi,r)$, corresponding to the azimuth, elevation, and radial distance, respectively. We assume the camera is always oriented to face the center of the object space.
To achieve camera-relative canonicalization, we rotate the normalized coordinates around the vertical axis (Z-axis) by the source camera’s azimuth $\theta$. This aligns the coordinate system with the viewpoint of the source camera. The transformation is applied around the center of the unit cube $\mathbf{h}$. We define the intermediate rotated coordinates $\mathbf{p}^{\prime\prime}$ as:
$\mathbf{p}^{\prime\prime}=R_{z}(\theta)(\mathbf{p}^{\prime}-\mathbf{h})+\mathbf{h},$ (4)
where $R_{z}(\theta)\in SO(3)$ is the standard 3D rotation matrix around the Z-axis:
$R_{z}(\theta)=\begin{pmatrix}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{pmatrix}$. (5)
This transformation rotates the coordinates in the X-Y plane (corresponding to the Red and Green channels in the pointmap) while preserving the Z coordinate (Blue channel). Notably, this definition is deliberately invariant to the camera elevation $\phi$ and distance $r$, a choice further elaborated in Appendix A.2.2.
Rescaling
Since the object was initially scaled to fit an axis-aligned unit cube (Eq. 3), the rotation in Eq. (4) may cause the coordinates $\mathbf{p}^{\prime\prime}$ to exceed the $0,1]^{3}$ bounds (illustrated in Appendix A.2.1). A final rescaling step is necessary to ensure the coordinates remain constrained within the unit cube.
To ensure the normalization is robust and predictable across different azimuth angles, we incorporate the theoretical bounds of the transformation. We define the theoretical range $L_{th}(\theta)$ and the theoretical minimum coordinate $m_{th}(\theta)$ achievable when rotating a unit cube by $\theta$ around its center:
$L_{th}(\theta)$ $=|\cos\theta|+|\sin\theta|,$ (6)
$m_{th}(\theta)$ $=0.5-\frac{1}{2}L_{th}(\theta).$ (7)
Let $\Omega^{\prime\prime}$ be the set of all rotated coordinates for the object. We calculate the observed minimum $m_{obs}$ and maximum $M_{obs}$ values across all points and all dimensions of $\Omega^{\prime\prime}$. The effective minimum shift $m_{eff}$ and the effective range $L_{eff}$ are calculated by clamping the observed values with the theoretical bounds. This ensures that the normalization is robust and consistent with the geometric constraints of the rotation:
$m_{eff}$ $=\max(m_{obs},m_{th}(\theta)),$ (8)
$L_{eff}$ $=\text{clip}(M_{obs}-m_{eff},\epsilon,L_{th}(\theta)),$ (9)
where $\epsilon$ is a small positive constant (e.g., $10^{-6}$) to prevent division by zero.
The final CROCS coordinates $\mathbf{p}_{\text{CROCS}}\in[0,1]^{3}$ are given by the uniformly rescaled coordinates:
$\mathbf{p}_{\text{CROCS}}=\frac{\mathbf{p}^{\prime\prime}-m_{eff}}{L_{eff}}.$ (10)
This procedure guarantees that the object remains tightly bounded by the unit cube in the camera-relative coordinate space.
A.2 Geometric Subtleties
A.2.1 CROCS Rescaling